# Privacy-Preserving Edge Federated Learning for Intelligent Mobile-Health Systems

### v1

@@ IV-B2 Performance and Overhead
IV-B3 Privacy and Overhead
In this section, we discuss the communication overhead, communication latency, and privacy of our SMC technique. We compare our techniques against three baseline SMC approaches, i.e., NOSMC, STSMC, and Shamir [114]. Table IV presents the analysis results for all the techniques. In this table, the latency/delay of transferring a packet from $P_{i}$ to $P_{j}$ is denoted by $L(P_{i},P_{j})$.
TABLE IV: Communication complexity, communication latency and privacy of SMC approaches ($n=$ number of data holders, $N=$ number of all parties including the mediator)
@@ IV-B3 Privacy and Overhead
    In regard to communication complexity, in the first round, all parties holding secret values ($n$ parties) send and receive $n-1$ messages. In the second round, $k-1$ parties send one message to $1$ party, and accordingly, that party receives $k-1$ messages. The total number of receiving and sending messages equals $2(n^{2}-n+k-1)$. Therefore, the communication overhead is of $O(n^{2})$. In regard to communication latency, in the first round, all parties holding secret values send other parties one message in parallel. In the second round, in the case where $k=n$ as it is the case in this paper, all parties holding secret values send one message to the mediator in parallel. The first and second rounds are performed sequentially. Therefore, the latency is equal to the delay of the last message received in the first round plus the delay of the last message received by the mediator in the second round. In regard to privacy, each party has $k$ private values, i.e., $k-1$ secret coefficients and $1$ secret value. If $k$ parties that received the initial results from one victim party collude in the first round, then, they can solve a system of $k$ equations ($k$ equations, $k$ unknowns) to reveal the victim party’s secret value.
@@
As the results in Table IV show, our proposed framework’s communication complexity is $O(n)$ and similar to the approach adopting no SMC scheme (NOSMC). This is particularly desirable for use when we are training a model on resource-constrained devices. If we had adopted Shamir’s scheme, at each round of secure aggregation, each edge device, at least, should have sent and received $n-1$ messages. This is while by adopting our secret sharing scheme, each edge device, at each round of secure aggregation, should merely send one message, which is considerably more efficient than Shamir’s technique. Our proposed framework’s communication latency is also similar to the approach adopting no SMC scheme (NOSMC). This is because both approaches have one round of message passing that is performed in parallel. Moreover, the privacy of our framework is similar to Shamir’s scheme, i.e., in both approaches, it takes at least $k$ data-holder parties to reveal a secret value.
In summary, the analysis results show that, in terms of communication complexity and communication latency, our proposed framework is as efficient as the NOSMC approach. At the same time, our framework protects the privacy of data holders as well as Shamir’s scheme.
IV-B4 Implementation on Amazon AWS Cloud
@@
Table V explains the three scenarios for our experiments in this paper. In all scenarios, the mediator is located in Sweden. In the first scenario, we have three data-holder parties located in England, France, and Germany. In the second scenario, we have five data-holder parties located in Canada, England, France, Germany, and India. In the third scenario, we have ten data-holder parties located in Canada, the United States, England, France, Germany, India, Singapore, South Korea, Japan, and Australia. The machines in all locations run Ubuntu 20.04 and include a 1-core 2.40 GHz CPU and 1 GB RAM, hence similar resources as existing mobile devices.
We measure the latency for one epoch of training a model based on our proposed framework for the aforementioned scenarios. The results of our experiments are reported in Table V. The reported results are the average results for 100 epochs of training. The latency of each epoch includes the latency of performing the required calculations on all involved machines, the mediator and data-holders, the latency of sending and receiving global parameters that the mediator shares with other parties, and the latency of communication for the secure aggregation process.
We performed two sets of experiments. In the first set of experiments, all data holders hold equal sizes of training data. As Table V shows, in this setting, the average latency for one epoch of training a model takes 0.3178 s, 0.6281 s, and 1.1428 s for Scenarios 1, 2, and 3, respectively. The results show how the distance between the mediator and data holders can affect the latency for training a model.
In the second set of experiments, we divided the training data among parties in each scenario. Therefore, in these experiments, parties in Scenario 1 hold larger training data compared to Scenario 2, and parties in Scenario 2 hold larger data compared to Scenario 3. As Table V shows, in Scenario 1, one epoch of training using our framework takes 1.4366 s on average. For Scenarios 2 and 3, one epoch of training takes 0.6726 s and 1.1428 s on average, respectively. In Scenario 1, due to the size of training data, which is larger than the size of training data in other scenarios, the latency is the highest on average. This is despite the fact that the geographical locations of data holders in this scenario are closer to the mediator compared to other scenarios. In Scenarios 2 and 3, the distance between data holders and the mediator increases. This leads to an increase in the latency for communication, and hence in Scenario 3, although each party holds smaller training data, the latency is higher than the latency in Scenario 2.
TABLE V: The scenarios for our experiments on Amazon’s AWS cloud and the latencies for one epoch training (average of 100 epochs)
@@ V Conclusions
In this work, we assume that the devices participating in the FL process have similar resources, e.g., e-Glass systems used by the patients are one single type of device. However, in the real-world IoT applications, the mobile devices are generally heterogeneous. The proposed scheme in this article may be extended to the settings with heterogeneous resource-constrained devices assuming synchronous updates to the federated model. In such settings, the latency of each round of federated updates among the heterogeneous resource-constrained devices will inevitably be determined by the slowest device participating in this round. In other words, while the proposed scheme can still be adopted in the context of heterogeneous devices, it has not been designed for such settings. In our future work, we will investigate privacy-preserving edge FL designed specifically for heterogeneous devices with heterogeneous computing power/resources.
Finally, in this work, we have mainly focused on the honest-but-curious privacy model. In our future work, we plan to extend this work towards other attack models, e.g., Byzantine, Poisoning, and Backdoor attacks [116, 117, 118, 119, 120].
Acknowledgment
This research has been partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), Swedish Research Council (VR), ELLIIT Strategic Research Environment, Swedish Foundation for Strategic Research (SSF), and European Union (EU) Interreg Program. The computations and data handling were enabled by resources provided by the Swedish National Infrastructure for Computing (SNIC) partially funded by the Swedish Research Council (VR) through grant agreement no. 2018-05973.

### v2

@@ 4.2.2 Performance and Overhead
4.2.3 Privacy and Overhead
In this section, we discuss the communication overhead, communication latency, and privacy of our SMC technique. We compare our techniques against three baseline SMC approaches, i.e., NOSMC, STSMC, and Shamir [114]. Table 4 presents the analysis results for all the techniques. In this table, the latency/delay of transferring a packet from $P_{i}$ to $P_{j}$ is denoted by $L(P_{i},P_{j})$.
Table 4: Communication complexity, communication latency and privacy of SMC approaches ($n=$ number of data holders, $N=$ number of all parties including the mediator)
@@ 4.2.3 Privacy and Overhead
    In regard to communication complexity, in the first round, all parties holding secret values ($n$ parties) send and receive $n-1$ messages. In the second round, $k-1$ parties send one message to $1$ party, and accordingly, that party receives $k-1$ messages. The total number of receiving and sending messages equals $2(n^{2}-n+k-1)$. Therefore, the communication overhead is of $O(n^{2})$. In regard to communication latency, in the first round, all parties holding secret values send other parties one message in parallel. In the second round, in the case where $k=n$, all parties holding secret values send one message to the mediator in parallel. The first and second rounds are performed sequentially. Therefore, the latency is equal to the delay of the last message received in the first round plus the delay of the last message received by the mediator in the second round. In regard to privacy, each party has $k$ private values, i.e., $k-1$ secret coefficients and $1$ secret value. If $k$ parties that received the initial results from one victim party collude in the first round, then, they can solve a system of $k$ equations ($k$ equations, $k$ unknowns) to reveal the victim party’s secret value.
@@
As the results in Table 4 show, our proposed framework’s communication complexity is $O(n)$ and similar to the NOSMC approach. This is particularly desirable for use when we are training a model on resource-constrained devices. If we had adopted Shamir’s scheme, at each round of secure aggregation, each edge device, at least, should have sent and received $n-1$ messages. This is while by adopting our secret sharing scheme, each edge device, at each round of secure aggregation, should merely send one message, which is considerably more efficient than Shamir’s technique. Our proposed framework’s communication latency is also similar to the NOSMC approach. This is because both approaches have one round of message passing that is performed in parallel. Moreover, the privacy of our framework is similar to Shamir’s scheme, i.e., in both approaches, it takes at least $k$ data-holder parties to reveal a secret value.
In summary, the analysis results show that, in terms of communication complexity and communication latency, our proposed framework is as efficient as the NOSMC approach. At the same time, our framework protects the privacy of data holders similar to Shamir’s scheme.
4.2.4 Implementation on Amazon AWS Cloud
@@
Table 5 explains the three scenarios for our experiments in this paper. In all scenarios, the mediator is located in Sweden. In the first scenario, we have three data-holder parties located in England, France, and Germany. In the second scenario, we have five data-holder parties located in Canada, England, France, Germany, and India. In the third scenario, we have ten data-holder parties located in Canada, the United States, England, France, Germany, India, Singapore, South Korea, Japan, and Australia. The machines in all locations run Ubuntu 20.04 and include a 1-core 2.40 GHz CPU and 1 GB RAM, hence similar resources as existing mobile devices.
We measure the latency for one epoch of training a model based on our proposed framework for the aforementioned scenarios. The results of our experiments are reported in Table 5. The reported results are the average results for 100 epochs of training. The latency of each epoch includes the latency of performing the required calculations on all involved machines, the mediator and data-holders, the latency of sending and receiving global parameters that the mediator shares with other parties, and the latency of communication for the secure aggregation process.
We performed two sets of experiments. In the first set of experiments, all data holders hold equal sizes of training data. As Table 5 shows, in this setting, the average latency for one epoch of training a model takes 0.3178 s, 0.6281 s, and 1.1428 s for Scenarios 1, 2, and 3, respectively. The results show how the distance between the mediator and data holders can affect the latency for training a model.
In the second set of experiments, we divided the training data among parties in each scenario. Therefore, in these experiments, parties in Scenario 1 hold more training data compared to Scenario 2, and parties in Scenario 2 hold more data compared to Scenario 3. As Table 5 shows, in Scenario 1, one epoch of training using our framework takes 1.4366 s on average. For Scenarios 2 and 3, one epoch of training takes 0.6726 s and 1.1428 s on average, respectively. In Scenario 1, due to the size of training data, which is larger than the size of training data in other scenarios, the latency is the highest on average. This is despite the fact that the geographical locations of data holders in this scenario are closer to the mediator compared to other scenarios. In Scenarios 2 and 3, the distance between data holders and the mediator increases. This leads to an increase in the latency for communication, and hence in Scenario 3, although each party holds less training data, the latency is higher than the latency in Scenario 2.
Table 5: The scenarios for our experiments on Amazon’s AWS cloud and the latencies for one epoch training (average of 100 epochs)
@@ 5 Conclusions
In this work, we assume that the devices participating in the FL process have similar resources, e.g., e-Glass systems used by the patients are one single type of device. However, in the real-world IoT applications, the mobile devices are generally heterogeneous. The proposed scheme in this article may be extended to the settings with heterogeneous resource-constrained devices assuming synchronous updates to the federated model. In such settings, the latency of each round of FL updates among the heterogeneous resource-constrained devices will inevitably be determined by the slowest device participating in this round. In other words, while the proposed scheme can still be adopted in the context of heterogeneous devices, it has not been designed for such settings. In our future work, we will investigate privacy-preserving edge FL designed specifically for heterogeneous devices with heterogeneous computing power/resources.
Finally, in this work, we have mainly focused on the honest-but-curious privacy model. In our future work, we plan to extend this work towards other attack models, e.g., poisoning, Byzantine, and backdoor attacks [116, 117, 118, 119, 120].